OpenStudy (anonymous):

simplify completely 12x+36/x^2-4x-21 and find the restrictions on the variable.

Parth (parthkohli):

\[\dfrac{12 x + 36}{x^2 - 4x - 21} = \dfrac{12x + 36}{x^2 - 7x + 3x - 21} = \dfrac{12x+ 36}{x(x - 7) + 3(x - 7)} = \dfrac{12x + 36}{(x +3)(x-7)}\]

OpenStudy (anonymous):

Ohhhh okay! I got a little confused. Thank you Parth :)

Parth (parthkohli):

Can you simplify it a little more?

OpenStudy (anonymous):

yes! I simplified it to 12/x-7 and the restrictions would be x does not equal 7 & x does not equal -3